OpenStudy (anonymous):

In 1977, the price of a scientific calculator was $175. In 2004, the price was $15. Assuming the change in price was linear, what was the approximate price of a scientific calculator in 1998?

OpenStudy (anonymous):

WHAT DID YOU DO TODAY?

OpenStudy (anonymous):

i'm at school (: what did you do today?

OpenStudy (anonymous):

I AM AT HOME DO YOU HAVE A DOG? I DO!!!

OpenStudy (anonymous):

no i dont, i want one. can you help with my problem? (:

OpenStudy (anonymous):

GO TO CCI.COM NOW

OpenStudy (anonymous):

y??????????????????????????????????????

OpenStudy (anonymous):

THAY HAVE GRAT DOGS!!!!

OpenStudy (anonymous):

I GOT MY DOG THARE!!!

OpenStudy (anonymous):

look, i need help with my math problem!!!!!, can you help or not (:

OpenStudy (anonymous):

in 27 years it dropped $160 so drops and average of \[\frac{160}{27}\] per year. in 21 year then it will drop \[21\times \frac{160}{27}=46\tfrac{2}{3}\]

OpenStudy (anonymous):

oops i made a mistake.